OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 4x − 3y = −12 and passes through the point (−3, 4).

OpenStudy (anonymous):

put the first equation into slop-intercept form

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

m= your slope. And for somehting to be parallel to it, it has to have the same slope

OpenStudy (anonymous):

\[3Y-4X=24\] as the slop equal to the slope of the given line =m ;then Y=mX=c

OpenStudy (anonymous):

y = 4/3x + 8 is that the answer?

OpenStudy (anonymous):

Help Im Lost!!

OpenStudy (anonymous):

Also This one: Write the equation of the line that is perpendicular to the line y = 2x + 2 and passes through the point (6, 3).

OpenStudy (anonymous):

Help me out please!

OpenStudy (anonymous):

the equation is 2Y+X+12

OpenStudy (anonymous):

ok a parallel line would have the same slope a perpindicular line would have a negative recipocal slope so start by determining the slope of the perpendicular line since the slope of the line is 2 the negative of 2 is -2 and the recipocal of -2 is -1/2 so the slope of the perpendicular is -1/2 since you have the slope, substitute the point and slope into the point slope formula to get your answer

OpenStudy (anonymous):

idk THanks though